# Diophantine approximation with sums of two squares

### v3

@@ 3. Proof of Corollary 1
We start the proof by the following lemma providing a bound for $r_{Q}(n)$.
@@
for some constant $\delta>0$ (see [8, Theorem 317]). It follows that
@@
$\alpha n=n_{1}+\frac{b}{q}+\frac{\theta n}{q^{2}},$ (3.4)
@@
$\alpha n||\ll_{Q}\frac{L}{q}+\frac{X}{q^{2}}=2q^{\beta-1}=2X^{-\frac{1-\beta}{1+\beta}}=2X^{-\gamma}\ll n^{-\gamma}.$ (3.5)
@@
$\{n\in\mathcal{A}_{Q}:n\leq 2X,\ ||n\alpha||<C_{1}n^{-\gamma}\}=:\mathcal{A}_{Q}(\alpha,X),\ \text{ say},$
@@
$\frac{XL}{q\log^{2}q}\ll S=\sum\limits_{\begin{subarray}{c}|b|\leq L\\ (b,q)=1\end{subarray}}\sum\limits_{\begin{subarray}{c}n\\ na\equiv b\bmod{q}\end{subarray}}r_{Q}(n)w(n)\ll X^{c/(\log\log X)}|A_{Q}(\alpha,X)|,$
@@
4. Reduction of Theorem 4
We detect the congruence condition on the right-hand side of (1.4) using the orthogonal property of additive characters, thus obtaining
$S=$ $\frac{1}{q}\sum_{\begin{subarray}{c}|b|\leq L\\ (b,q)=1\end{subarray}}\sum_{n}r_{Q}(n)w(n)\sum_{g\bmod q}e_{q}\left(n-\bar{a}b)g\right)$ (4.1)
$=$ $\frac{1}{q}\sum_{\begin{subarray}{c}|b|\leq L\\ (b,q)=1\end{subarray}}\sum_{n}r_{Q}(n)w(n)\sum_{k|q}\ \sideset{}{{}^{*}}{\sum}_{h\bmod k}e_{k}\left(n-\bar{a}b)h\right),$ (4.2)
where $a\bar{a}\equiv 1\bmod q$ and the asterisk above indicates that $(h,k)=1$. As the weight function $w(n)$ is compactly supported, all the summations in $S$ run over finite ranges and so we can interchange the order of the summations. Hence,
$S=\frac{1}{q}\sum_{\begin{subarray}{c}|b|\leq L\\ (b,q)=1\end{subarray}}\sum_{k|q}\ \sideset{}{{}^{*}}{\sum}_{h\bmod k}e_{k}(-\bar{a}bh)\sum_{n}r_{Q}(n)w(n)e_{k}(nh).$ (4.3)
Now applying Lemma 3 to the inner-most sum, we have
$S$ $=\frac{2\pi}{q\Delta^{1/2}}\left(\int\limits_{0}^{\infty}w(x)dx\right)\sum_{\begin{subarray}{c}|b|\leq L\\ (b,q)=1\end{subarray}}\sum_{k|q}\frac{1}{k^{2}}\sideset{}{{}^{*}}{\sum}_{h\bmod k}G_{Q}(k,h)e_{k}(-\bar{a}bh)$
$+\frac{2\pi}{q\Delta^{1/2}}\hskip-5.69046pt\sum_{\begin{subarray}{c}|b|\leq L\\ (b,q)=1\end{subarray}}\hskip-5.69046pt\sum_{k|q}\ \sideset{}{{}^{*}}{\sum}_{h\bmod k}\hskip-8.5359pte_{k}(-\bar{a}bh)\sum_{n}\tilde{r}_{Q^{*}}(n,k,h)e_{k_{1}}(-\bar{h}\bar{\Delta}_{0}\delta_{1}^{-1}n)\tilde{w}_{k}(n)$
$=:T_{1}+T_{2}\ \text{(}say),$ (4.4)
@@ 4. Reduction of Theorem 4
$\tilde{w}_{k}(n):=\frac{1}{k}\int\limits_{0}^{\infty}J_{0}\left(4\pi\frac{\sqrt{nx/\Delta_{0}}}{k}\right)w(x)dx.$ (4.5)
Recall (2.12), which says that
$\tilde{r}_{Q^{*}}(n,k,h)=k^{-1}\sum_{\begin{subarray}{c}Q^{*}(\vec{x})=n\end{subarray}}\ \sum_{\vec{y}\bmod k}e_{k}(hQ(\vec{y}))e_{\delta_{1}}(\vec{y}\cdot gVB\vec{x}),$ (4.6)
where $\delta_{0}=\delta_{0}(k),\delta_{1}=\delta_{1}(k),g$ are given in (2.1), (2.2), (2.5), respectively and $Q^{*}$ is the quadratic form associated to the matrix (2.11). Recall also our assumption $(2\Delta,q)=1$. As $k|q$, it follows that $\delta=\delta_{0}\delta_{1}=(\Delta,k)=1$, i.e., $\delta_{0}=\delta_{1}=1$. Thus, the matrices $B,V$ have integer entries. Since, $\vec{x},\vec{y}$ are integer vectors, we notice that $e_{\delta_{1}}(\vec{y}\cdot gVB\vec{x})=1$. Therefore, by (4.6), (2.3) and (2.8), we have
$\tilde{r}_{Q^{*}}(n,k,h)$ $=k^{-1}\sum_{\begin{subarray}{c}Q^{*}(\vec{x})=n\end{subarray}}\ \sum_{\vec{y}\bmod k}e_{k}(hQ(\vec{y}))$
$=k^{-1}\sum_{\begin{subarray}{c}Q^{*}(\vec{x})=n\end{subarray}}G_{Q}(k,h)$
$=k^{-1}\sum_{\begin{subarray}{c}Q^{*}(\vec{x})=n\end{subarray}}\left(\frac{h}{N(k,Q)}\right)G_{Q}(k,1)$
$=k^{-1}G_{Q}(k,1)r_{Q*}(n)\left(\frac{h}{N(k,Q)}\right).$ (4.7)
@@
Let $q$ be given as in (3.1) and $N(q,Q)$ be as in (2.9). Assume that $(2\Delta,q)=1$. Then for $k|q$, we have
@@
$A\vec{x}\equiv 0\bmod{k}\Longleftrightarrow\begin{cases}2a_{1}x+b_{1}y\equiv 0\bmod{k}\\ b_{1}x+2c_{1}y\equiv 0\bmod{k},\end{cases}$ (4.9)
where $\vec{x}=(x,y)$. Observe that the determinant of the coefficient matrix
@@
By the above Lemma 7, we have
$\tilde{r}_{Q^{*}}(n,k,h)=k^{-1}G_{Q}(k,1)r_{Q*}(n)$ (4.10)
from (4).
@@ 5. Evaluation of the main term $T_{1}$
Observe that $q$ is odd since we assumed that $(2\Delta,q)=1$. Since $k|q$, by Lemma 1(i) and Lemma 7, we have
@@
Therefore, the main term $T_{1}$ in (4) turns into
$T_{1}$ $=\frac{2\pi}{q\Delta^{1/2}}\left(\int\limits_{0}^{\infty}w(x)dx\right)\sum_{\begin{subarray}{c}|b|\leq L\\ (b,q)=1\end{subarray}}\sum_{k|q}\frac{1}{k^{2}}G_{Q}(k,1)\sideset{}{{}^{*}}{\sum}_{h\bmod k}e_{k}(-\bar{a}bh)$
$=\frac{2\pi}{q\Delta^{1/2}}\left(\int\limits_{0}^{\infty}w(x)dx\right)\sum_{\begin{subarray}{c}|b|\leq L\\ (b,q)=1\end{subarray}}\sum_{k|q}\frac{\mu(k)}{k^{2}}G_{Q}(k,1)$
$=\frac{2\pi}{q\Delta^{1/2}}\left(\int\limits_{0}^{\infty}w(x)dx\right)\left(\frac{\varphi(q)}{q}\cdot 2L+O(d(q))\right)\sum_{k|q}\frac{\mu(k)}{k^{2}}G_{Q}(k,1),$ (5.2)
where $d(q)$ denotes the number of divisors of $q$. From the first line to the second line, we used the fact that the Ramanujan sum
$c_{k}(\bar{a}b)=\sideset{}{{}^{*}}{\sum}_{h\bmod k}e_{k}(-\bar{a}bh)$ (5.3)
satisfies $c_{k}(\bar{a}b)=\mu(k)$ if $(\bar{a}b,k)=1$.
@@
where the products are restricted to prime divisors $p$ of $q$. This, together with (5) gives
$T_{1}=\frac{4\pi\varphi(q)L}{\Delta^{1/2}q^{2}}\prod_{\begin{subarray}{c}p|q\\ p\text{ prime }\end{subarray}}\left(1-\frac{G_{Q}(p,1)}{p^{2}}\right)\int\limits_{0}^{\infty}w(x)dx+O_{\varepsilon}\left(\frac{Xq^{\varepsilon}}{q}\right).$ (5.6)
Therefore, using Mertens formula, from (5.5), (5.6) and
$\frac{\varphi(q)}{q}=\prod\limits_{p|q}\left(1-\frac{1}{p}\right),$
we deduce the following lemma.
@@
$\frac{XL}{q(\log q)^{2}}\ll T_{1}\ll\frac{XL}{q}.$
@@ 6. Evaluation of $T_{2}$
Inserting the formula (4) of $\tilde{r}_{Q^{*}}(n,k,h)$ in (4), we have
$T_{2}$ $=\frac{2\pi}{q\Delta^{1/2}}\sum_{\begin{subarray}{c}b\leq L\\ (b,q)=1\end{subarray}}\sum_{k|q}k^{-1}G_{Q}(k,1)\sum_{n}r_{Q^{*}}(n)\tilde{w}_{k}(n)\sideset{}{{}^{*}}{\sum}_{h\bmod k}e_{k}(-\bar{a}bh-\bar{\Delta}_{0}\bar{h}n)$
$=\frac{2\pi}{q\Delta^{1/2}}\sum_{k|q}\frac{G_{Q}(k,1)}{k}\sum_{n}\sum_{\begin{subarray}{c}b\leq L\\ (b,q)=1\end{subarray}}r_{Q^{*}}(n)\tilde{w}_{k}(n)S(\bar{\Delta}_{0}n,\bar{a}b;k)$
$=\frac{2\pi}{q\Delta^{1/2}}\sum_{k|q}\frac{G_{Q}(k,1)}{k}\sum_{n}\sum_{\begin{subarray}{c}b\leq L\\ (b,q)=1\end{subarray}}r_{Q^{*}}(n)\tilde{w}_{k}(n)S(n,\bar{\Delta}_{0}\bar{a}b;k),$ (6.1)
since $(\bar{\Delta}_{0},k)=1$.
In order to bound $\tilde{w}_{k}(n)$, we use the following facts about the Bessel functions, which are available in many places (see [2, equations (9) and (10)]). For any $\beta>0$ and positive integer $j$, by repeated integrations by parts, we have
@@
Also, for any non negative integer $j$, we have
$J_{j}(x)\ll_{\varepsilon}\frac{1}{\sqrt{1+x}}+\frac{1}{x^{j-\varepsilon}}$ (6.3)
for arbitrarily small $\varepsilon>0$. Using the trivial bound $J_{j}(x)\ll_{\varepsilon}x^{\varepsilon}$ together with the fact that $w^{(j)}(x)\ll x^{-j}$, and recalling that $w(x)$ has compact support in $[X,2X]$, we infer from (4.5) and (6.2) that
$\tilde{w}_{k}(n)$ $\ll_{j,\varepsilon}\frac{1}{k}\left(\frac{k}{\sqrt{n}}\right)^{j}\int\limits_{X}^{2X}x^{-j}x^{j/2}x^{\varepsilon}dx$
$\ll_{j,\varepsilon}\frac{1}{k}\left(\frac{k}{\sqrt{n}}\right)^{j}X^{-j/2+1+\varepsilon}$
$=\frac{X^{1+\varepsilon}}{k}\left(\frac{k}{\sqrt{Xn}}\right)^{j}$ (6.4)
for any $j\geq 1$. Thus, taking $j$ sufficiently large, we see that $\tilde{w}_{k}(n)$ is negligible (that is, say $\tilde{w}_{k}(n)\ll X^{-20250}$) unless
$\frac{k}{\sqrt{Xn}}\geq X^{-\varepsilon/2},\text{ that is, }n\leq\frac{k^{2}}{X^{1-\varepsilon}}.$ (6.5)
Using this, it follows from (6) that
$T_{2}$ $=\frac{2\pi}{q\Delta^{1/2}}\sum_{k|q}\frac{G_{Q}(k,1)}{k}\sum_{n\leq k^{2}/X^{1-\varepsilon}}\sum_{\begin{subarray}{c}b\leq L\\ (b,q)=1\end{subarray}}r_{Q^{*}}(n)\tilde{w}_{k}(n)S(n,\bar{\Delta}_{0}\bar{a}b;k)+O(X^{-2025}).$ (6.6)
@@
Set $\gamma(n):=r_{Q^{*}}(n)\tilde{w}_{k}(n)$ if $1\leq n\leq k^{2}/X^{1-\varepsilon}$ and $\gamma(n):=0$ otherwise, where $r_{Q^{*}}(n),\tilde{w}_{k}(n)$ are given as in (6.6). Then we have
[…]
@@
Taking $j=0$ in (6), we have
$|\tilde{w}_{k}(n)|\leq\frac{X^{1+\varepsilon}}{k}.$ (6.8)
@@
Let $T_{2}$ be given in (6). Then we have
$T_{2}\ll_{\varepsilon}q^{\varepsilon}(X^{1/2}L^{1/4}+X^{1/2}q^{-1/2}L+q^{1/2}L^{1/2})$ (6.10)
@@
In order to apply Lemma 4, we need to settle the coprimality condition $(b,q)=1$, in the inner sum. We do this by Möbius inversion. We have
$\sum_{n\leq k^{2}/X^{1-\varepsilon}}\sum_{\begin{subarray}{c}b\leq L\\ (b,q)=1\end{subarray}}r_{Q^{*}}(n)\tilde{w}_{k}(n)S(n,\bar{\Delta}_{0}\bar{a}b;k)$
$=$ $\sum_{n\leq k^{2}/X^{1-\varepsilon}}\sum_{\begin{subarray}{c}b\leq L\end{subarray}}\left(\sum_{d|(b,q)}\mu(d)\right)r_{Q^{*}}(n)\tilde{w}_{k}(n)S(n,\bar{\Delta}_{0}\bar{a}b;k)$
$=$ $\sum_{d|q}\mu(d)\sum_{n\leq k^{2}/X^{1-\varepsilon}}\sum_{\begin{subarray}{c}b\leq L/d\end{subarray}}r_{Q^{*}}(n)\tilde{w}_{k}(n)S(n,\bar{\Delta}_{0}\bar{a}db;k).$
Now $(\bar{\Delta}_{0}\bar{a}d,k)=(d,k)=s,$ say. Applying the second bound of Lemma 4 with $\gamma(n)=r_{Q^{*}}(n)\tilde{w}_{k}(n)$ to
$\sum_{n\leq k^{2}/X^{1-\varepsilon}}\sum_{b\leq L/d}r_{Q^{*}}(n)\tilde{w}_{k}(n)S(n,\bar{\Delta}_{0}\bar{a}db;k),$
we obtain
$\sum_{n\leq k^{2}/X^{1-\varepsilon}}\sum_{\begin{subarray}{c}b\leq L/d\end{subarray}}r_{Q^{*}}(n)\tilde{w}_{k}(n)S(n,\bar{\Delta}_{0}\bar{a}db;k)$
[…]
@@
where the last line is due to the bound of $\gamma||$ from Lemma 10. Therefore,
$\sum_{d|k}\mu(d)\sum_{n\leq k^{2}/X^{1-\varepsilon}}\sum_{\begin{subarray}{c}b\leq L/d\end{subarray}}r_{Q^{*}}(n)\tilde{w}_{k}(n)S(n,\bar{\Delta}_{0}\bar{a}db;k)$
$\ll$ $(kX)^{3\varepsilon}k^{\varepsilon/2}\left\{X^{1/2}kL^{1/4}+X^{1/2}k^{1/2}L+k^{3/2}L^{1/2}\right\}$

### v4

@@ 3. Proof of Corollary 1
We start the proof of Corollary 1 by the following lemma providing a bound for $r_{Q}(n)$.
@@
for some constant $\delta>0$ (see [7, Theorem 317]). It follows that
@@
$\alpha n=n_{1}+\frac{b}{q}+\frac{\theta n}{q^{2}},$
@@
$\|\alpha n\|\ll_{Q}\frac{L}{q}+\frac{X}{q^{2}}=2q^{\beta-1}=2X^{-(1-\beta)/(1+\beta)}=2X^{-\gamma}\ll n^{-\gamma}.$
@@
$\{n\in\mathcal{A}_{Q}:n\leq 2X,\ \|n\alpha\|<C_{1}n^{-\gamma}\}=:\mathcal{A}_{Q}(\alpha,X),\ \text{ say},$
@@
$\frac{XL}{q\log^{2}q}\ll S=\sum\limits_{\begin{subarray}{c}|b|\leq L\\ (b,q)=1\end{subarray}}\sum\limits_{\begin{subarray}{c}n\in\mathbb{Z}\\ na\equiv b\bmod{q}\end{subarray}}r_{Q}(n)w(n)\ll X^{c/(\log\log X)}|A_{Q}(\alpha,X)|,$
@@
4. Application of Voronoi summation
Our proof of Theorem 4 begins with an application of Voronoi summation. We first detect the congruence condition on the right-hand side of (1.3) using the orthogonal property of additive characters, thus obtaining
$S=$ $\frac{1}{q}\sum_{\begin{subarray}{c}|b|\leq L\\ (b,q)=1\end{subarray}}\sum_{n\in\mathbb{Z}}r_{Q}(n)w(n)\sum_{g\bmod q}e_{q}\left((n-\overline{a}b)g\right)$
$=$ $\frac{1}{q}\sum_{\begin{subarray}{c}|b|\leq L\\ (b,q)=1\end{subarray}}\sum_{n\in\mathbb{Z}}r_{Q}(n)w(n)\sum_{k|q}\ \sideset{}{{}^{*}}{\sum}_{h\bmod k}e_{k}\left((n-\overline{a}b)h\right),$
where $a\overline{a}\equiv 1\bmod q$, and the asterisk above indicates that $(h,k)=1$. As the weight function $w(n)$ is compactly supported, all the summations in $S$ run over finite ranges and so we can interchange the order of summations. Hence,
$S=\frac{1}{q}\sum_{\begin{subarray}{c}|b|\leq L\\ (b,q)=1\end{subarray}}\sum_{k|q}\ \sideset{}{{}^{*}}{\sum}_{h\bmod k}e_{k}(-\overline{a}bh)\sum_{n\in\mathbb{Z}}r_{Q}(n)w(n)e_{k}(nh).$
Now applying Lemma 3 to the inner-most sum, we get
$\begin{split}S=&\frac{2\pi}{q\Delta^{1/2}}\left(\int\limits_{0}^{\infty}w(x)dx\right)\sum_{\begin{subarray}{c}|b|\leq L\\ (b,q)=1\end{subarray}}\sum_{k|q}\frac{1}{k^{2}}\sideset{}{{}^{*}}{\sum}_{h\bmod k}G_{Q}(k,h)e_{k}(-\overline{a}bh)+\\ &\frac{2\pi}{q\Delta^{1/2}}\sum_{\begin{subarray}{c}|b|\leq L\\ (b,q)=1\end{subarray}}\sum_{k|q}\ \sideset{}{{}^{*}}{\sum}_{h\bmod k}\hskip-8.5359pte_{k}(-\overline{a}bh)\sum_{n\in\mathbb{Z}}\tilde{r}_{Q^{*}}(n,k,h)e_{k}(-\overline{h}\overline{\Delta}n)\tilde{w}_{k}(n)\\ =:&T_{1}+T_{2}\ \text{(}say),\end{split}$ (4.1)
@@ 4. Application of Voronoi summation
$\tilde{w}_{k}(n):=\frac{1}{k}\int\limits_{0}^{\infty}J_{0}\left(4\pi\frac{\sqrt{nx/\Delta}}{k}\right)w(x)dx.$ (4.2)
Recalling our assumption $(2\Delta h,q)=1$, and using (2.1), (2.2) and (2.4), we obtain
$\begin{split}\tilde{r}_{Q^{*}}(n,k,h)&=k^{-1}\sum_{\begin{subarray}{c}Q^{*}(\vec{x})=n\end{subarray}}\ \sum_{\vec{y}\bmod k}e_{k}(hQ(\vec{y}))\\ &=k^{-1}\sum_{\begin{subarray}{c}Q^{*}(\vec{x})=n\end{subarray}}G_{Q}(k,h)\\ &=k^{-1}\sum_{\begin{subarray}{c}Q^{*}(\vec{x})=n\end{subarray}}\left(\frac{h}{N(k,Q)}\right)G_{Q}(k,1)\\ &=k^{-1}G_{Q}(k,1)r_{Q*}(n)\left(\frac{h}{N(k,Q)}\right).\end{split}$ (4.3)
@@
Assume that $(2\Delta,q)=1$. Then for $k|q$, we have
@@
$A\vec{x}\equiv 0\bmod{k}\Longleftrightarrow\begin{cases}2a_{1}x+b_{1}y\equiv 0\bmod{k}\\ b_{1}x+2c_{1}y\equiv 0\bmod{k},\end{cases}$
where $\vec{x}=(x,y)^{T}$. Observe that the determinant of the coefficient matrix
@@
Combining the above Lemma 7 with (4.3), we see that
$\tilde{r}_{Q^{*}}(n,k,h)=k^{-1}G_{Q}(k,1)r_{Q*}(n).$ (4.5)
@@ 5. Evaluation of the main term $T_{1}$
Using $k|q$ and Lemmas 1 and 7, we have
@@
Therefore, the main term $T_{1}$ in (4.1) turns into
$\begin{split}T_{1}=&\frac{2\pi}{q\Delta^{1/2}}\left(\int\limits_{0}^{\infty}w(x)dx\right)\sum_{\begin{subarray}{c}|b|\leq L\\ (b,q)=1\end{subarray}}\sum_{k|q}\frac{1}{k^{2}}G_{Q}(k,1)\sideset{}{{}^{*}}{\sum}_{h\bmod k}e_{k}(-\overline{a}bh)\\ =&\frac{2\pi}{q\Delta^{1/2}}\left(\int\limits_{0}^{\infty}w(x)dx\right)\sum_{\begin{subarray}{c}|b|\leq L\\ (b,q)=1\end{subarray}}\sum_{k|q}\frac{\mu(k)}{k^{2}}G_{Q}(k,1)\\ =&\frac{2\pi}{q\Delta^{1/2}}\left(\int\limits_{0}^{\infty}w(x)dx\right)\left(\frac{\varphi(q)}{q}\cdot 2L+O(\tau(q))\right)\sum_{k|q}\frac{\mu(k)}{k^{2}}G_{Q}(k,1),\end{split}$ (5.2)
where $\tau(q)$ denotes the number of divisors of $q$. Here we have used the fact that the Ramanujan sum
$c_{k}(\overline{a}b)=\sideset{}{{}^{*}}{\sum}_{h\bmod k}e_{k}(-\overline{a}bh)$
satisfies $c_{k}(\overline{a}b)=\mu(k)$ if $(\overline{a}b,k)=1$.
@@
where the products are restricted to prime divisors $p$ of $q$. This, together with (5.2) gives
$T_{1}=\frac{4\pi\varphi(q)L}{\Delta^{1/2}q^{2}}\prod_{\begin{subarray}{c}p|q\\ p\text{ prime }\end{subarray}}\left(1-\frac{G_{Q}(p,1)}{p^{2}}\right)\int\limits_{0}^{\infty}w(x)dx+O_{\varepsilon}\left(\frac{Xq^{\varepsilon}}{q}\right),$ (5.5)
where we have used the divisor bound $\tau(q)\ll_{\varepsilon}q^{\varepsilon}$. Taking the well-known bound
$\frac{\varphi(q)}{q}=\prod\limits_{p|q}\left(1-\frac{1}{p}\right)\gg\frac{1}{\log\log q},$
into account, we deduce the following lemma from (5.4) and (5.5).
@@
$\frac{q^{2\beta}}{(\log\log q)^{2}}=\frac{XL}{q(\log\log q)^{2}}\ll T_{1}\ll\frac{XL}{q}=q^{2\beta}.$
@@ 6. Evaluation of $T_{2}$
Inserting the formula (4.5) for $\tilde{r}_{Q^{*}}(n,k,h)$ into (4.1), we obtain
$\begin{split}T_{2}&=\frac{2\pi}{q\Delta^{1/2}}\sum_{\begin{subarray}{c}b\leq L\\ (b,q)=1\end{subarray}}\sum_{k|q}k^{-1}G_{Q}(k,1)\sum_{n\in\mathbb{Z}}r_{Q^{*}}(n)\tilde{w}_{k}(n)\sideset{}{{}^{*}}{\sum}_{h\bmod k}e_{k}(-\overline{a}bh-\overline{\Delta}\overline{h}n)\\ &=\frac{2\pi}{q\Delta^{1/2}}\sum_{k|q}\frac{G_{Q}(k,1)}{k}\sum_{n\in\mathbb{Z}}\sum_{\begin{subarray}{c}b\leq L\\ (b,q)=1\end{subarray}}r_{Q^{*}}(n)\tilde{w}_{k}(n)S(\overline{\Delta}n,\overline{a}b;k)\\ &=\frac{2\pi}{q\Delta^{1/2}}\sum_{k|q}\frac{G_{Q}(k,1)}{k}\sum_{n\in\mathbb{Z}}\sum_{\begin{subarray}{c}b\leq L\\ (b,q)=1\end{subarray}}r_{Q^{*}}(n)\tilde{w}_{k}(n)S(n,\overline{\Delta}\overline{a}b;k),\end{split}$ (6.1)
since $(\overline{\Delta},k)=1$.
In order to bound $\tilde{w}_{k}(n)$, we use the following facts about Bessel functions, which are available at many places (see [2, equations (9) and (10)], for example). For any $\beta>0$ and positive integer $j$, by repeated integrations by parts, we have
@@
Also, for any non-negative integer $j$ and arbitrarily small $\varepsilon>0$, we have
$J_{j}(x)\ll_{\varepsilon}\frac{1}{\sqrt{1+x}}+\frac{1}{x^{j-\varepsilon}}.$ (6.3)
Using the trivial bound $J_{j}(x)\ll_{\varepsilon}x^{\varepsilon}$ together with the fact that $w^{(j)}(x)\ll x^{-j}$, and recalling that $w(x)$ has compact support in $[X,2X]$, we infer from (4.2) and (6.2) that
$\begin{split}\tilde{w}_{k}(n)&\ll_{j,\varepsilon}\frac{1}{k}\left(\frac{k}{\sqrt{n}}\right)^{j}\int\limits_{X}^{2X}x^{-j}x^{j/2}x^{\varepsilon}dx\\ &\ll_{j,\varepsilon}\frac{1}{k}\left(\frac{k}{\sqrt{n}}\right)^{j}X^{-j/2+1+\varepsilon}\\ &=\frac{X^{1+\varepsilon}}{k}\left(\frac{k}{\sqrt{Xn}}\right)^{j}\end{split}$ (6.4)
for any $j\geq 1$. Thus, taking $j$ sufficiently large, we see that $\tilde{w}_{k}(n)$ is negligible (that is, $\tilde{w}_{k}(n)\ll X^{-20260}$) unless
$\frac{k}{\sqrt{Xn}}\geq X^{-\varepsilon/2},\text{ that is, }n\leq\frac{k^{2}}{X^{1-\varepsilon}}.$
Hence, it follows from (6.1) that
$T_{2}$ $=\frac{2\pi}{q\Delta^{1/2}}\sum_{k|q}\frac{G_{Q}(k,1)}{k}\sum_{n\leq k^{2}/X^{1-\varepsilon}}\sum_{\begin{subarray}{c}b\leq L\\ (b,q)=1\end{subarray}}r_{Q^{*}}(n)\tilde{w}_{k}(n)S(n,\overline{\Delta}\overline{a}b;k)+O(X^{-2026}).$ (6.5)
@@
Set
$\gamma(n):=\begin{cases}r_{Q^{*}}(n)\tilde{w}_{k}(n)&\mbox{ if }1\leq n\leq k^{2}/X^{1-\varepsilon},\\ 0&\mbox{ otherwise,}\end{cases}$
where $r_{Q^{*}}(n)$ and $\tilde{w}_{k}(n)$ are given as in (6.5). Then we have
$\|\bm{\gamma}\|_{2}:=\left(\sum_{n\in\mathbb{Z}}|\gamma(n)|^{2}\right)^{1/2}\ll_{\varepsilon}(kX)^{2\varepsilon}X^{1/2}.$
@@
Taking $j=0$ in (6.4), we get
$|\tilde{w}_{k}(n)|\leq\frac{X^{1+\varepsilon}}{k}.$
@@
Let $T_{2}$ be as given in (6.1). Then we have
$T_{2}\ll_{\varepsilon}q^{\varepsilon}(X^{1/2}L^{1/4}+X^{1/2}q^{-1/2}L+q^{1/2}L^{1/2})$
@@
In order to apply Lemma 4, we need to remove the coprimality condition $(b,q)=1$ in the innermost sum. We do this by Möbius inversion, obtaining
$\sum_{n\leq k^{2}/X^{1-\varepsilon}}\sum_{\begin{subarray}{c}b\leq L\\ (b,q)=1\end{subarray}}r_{Q^{*}}(n)\tilde{w}_{k}(n)S(n,\overline{\Delta}\overline{a}b;k)$
$=$ $\sum_{n\leq k^{2}/X^{1-\varepsilon}}\sum_{\begin{subarray}{c}b\leq L\end{subarray}}\left(\sum_{d|(b,q)}\mu(d)\right)r_{Q^{*}}(n)\tilde{w}_{k}(n)S(n,\overline{\Delta}\overline{a}b;k)$
$=$ $\sum_{d|q}\mu(d)\sum_{n\leq k^{2}/X^{1-\varepsilon}}\sum_{\begin{subarray}{c}b\leq L/d\end{subarray}}r_{Q^{*}}(n)\tilde{w}_{k}(n)S(n,\overline{\Delta}\overline{a}db;k).$
Set $s:=(\overline{\Delta}\overline{a}d,k)=(d,k)$. Then applying Lemma 4 with $\gamma(n)=r_{Q^{*}}(n)\tilde{w}_{k}(n)$ to the bilinear sum over $n$ and $b$ in the last line and choosing $S_{1}(M,N,q,d)$ as in (2.6), we obtain
$\sum_{n\leq k^{2}/X^{1-\varepsilon}}\sum_{\begin{subarray}{c}b\leq L/d\end{subarray}}r_{Q^{*}}(n)\tilde{w}_{k}(n)S(n,\overline{\Delta}\overline{a}db;k)$
$\ll$ $X^{\varepsilon}\|\bm{\gamma}\|_{2}(k^{2}/X)^{1/2}(L/d)k^{1/2+o(1)}\left\{(k^{2}/X)^{-1/2}\left((L/d)^{-3/4}k^{1/2}+s^{1/2}\right)+(L/d)^{-1/2}\right\}$
$\ll$ $(kX)^{\varepsilon}\|\bm{\gamma}\|_{2}\left\{kL^{1/4}d^{-1/4}+k^{1/2}Ld^{-1/2}+k^{3/2}L^{1/2}d^{-1/2}X^{-1/2}\right\}$
@@
where the last line is due to the bound for $\|\bm{\gamma}\|_{2}$ in Lemma 10. Therefore,
$\sum_{d|k}\mu(d)\sum_{n\leq k^{2}/X^{1-\varepsilon}}\sum_{\begin{subarray}{c}b\leq L/d\end{subarray}}r_{Q^{*}}(n)\tilde{w}_{k}(n)S(n,\overline{\Delta}\overline{a}db;k)$
$\ll$ $(kX)^{3\varepsilon}k^{\varepsilon/2}\left\{X^{1/2}kL^{1/4}+X^{1/2}k^{1/2}L+k^{3/2}L^{1/2}\right\}.$